find the equation of the line that contains (0,4) and parallel to y = 1/2x +2
part two
find the equation of the line that contains (0,4) and perpendicular to y = 1/2x +2

Answer by stanbon(75887)

Find the equation of the line that contains (0,4) and parallel to y = 1/2x +2
---
slope = 1/2
intercept = 4
-------------------
Equation: y = (1/2)x + 4
=====================================
part two
Find the equation of the line that contains (0,4)
and perpendicular to y = 1/2x +2
---
slope = -2
intercept = 4
-------------------
Equation: Y = -2x+4
=========================
Cheers,
Stan H.